OpenStudy (jessiegonzales):

HELP!!!! MEDAL!! ANYONE!!!! 5. m

OpenStudy (jessiegonzales):

@happytales @kropot72

OpenStudy (campbell_st):

well by bisecting <ABC the angles formed <ABD and <DBC are equal so equate the equations 5x = 3x + 10 solve for x then you will be able to find the measure of angles or value of x

OpenStudy (jessiegonzales):

is it 5?

OpenStudy (campbell_st):

well x = 5 I don't know if you just needed to find x or find the angle measure...

OpenStudy (jessiegonzales):

it says find m<ABC so thats 5? @campbell_st

OpenStudy (campbell_st):

ok... so if x =5 then you have the 2 equation 5x and 3x + 10 substitute and then add the angle measures together to get the measure of <ABC

OpenStudy (campbell_st):

the 2 smaller angles must add to the larger angle...

OpenStudy (jessiegonzales):

so the two smaller angles combines makes 30* right? @campbell_st

OpenStudy (campbell_st):

well not quite 5 * 5 = 25 and 3*5 + 10 = 25 so the measure of <ABC = 25 + 25

OpenStudy (jessiegonzales):

Thanks so much so the answer is 50?

OpenStudy (campbell_st):

yes